OpenStudy (anonymous):

distance between (5,-3), (7,1)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[\sqrt{20} =2\sqrt{5}\]

OpenStudy (anonymous):

\[d=\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\]

OpenStudy (anonymous):

\[d=\sqrt{(7-5)^{2}+(1-(-3))^{2}}\]

OpenStudy (anonymous):

\[d=\sqrt{(2)^{2}+(4)^{2}}\]

OpenStudy (anonymous):

\[d=\sqrt{(4+16)}\]

OpenStudy (anonymous):

\[d=\sqrt{20}\]

OpenStudy (anonymous):

\[\sqrt{20}=\sqrt{4}x \sqrt{5}\]

OpenStudy (anonymous):

\[\sqrt{4}x \sqrt{5}=2\sqrt{5}\] so I guess you could say I agree with Eliza27